# Optimal control of the future via prospective learning with control

### v2

@@
Optimal control of the future is the next frontier for AI. Current approaches to this problem are typically rooted in either reinforcement learning (RL). While powerful, this learning framework is mathematically distinct from supervised learning, which has been the main workhorse for the recent achievements in AI. Moreover, RL typically operates in a stationary environment with episodic resets, limiting its utility to more realistic settings. Here, we extend supervised learning to address learning to control in non-stationary, reset-free environments. Using this framework, called ”Prospective Learning with Control (PL+C)”, we prove that under certain fairly general assumptions, empirical risk minimization (ERM) asymptotically achieves the Bayes optimal policy. We then consider a specific instance of prospective learning with control, foraging—which is a canonical task for any mobile agent—be it natural or artificial. We illustrate that modern RL algorithms fail to learn in these non-stationary reset-free environments, and even with modifications, they are orders of magnitude less efficient than our prospective foraging agents.
@@ Introduction
9:Online updates $\hat{g}^{i}$ and $\hat{g}^{p}$ and $x_{s}$
10:for $s=1>T$
11:  $\mathcal{C}\leftarrow\textsc{AllPaths}(x_{s},H)$ $\triangleright$ each $x^{(s:H+s)}\in\mathcal{C}$ is a length-$H$ trajectory, $(x^{(s)},\dots,x^{(H+s)})$
12:  for $x^{(s:H+s)}\in\mathcal{C}$
13:  $\hat{G}(x^{(s:H+s})\leftarrow\sum_{h=1}^{H}\gamma^{\,h-1}\;\mathrm{\hat{g}^{i}}(\varphi(x^{(h)},s{+}h))$ $\triangleright$ approx. finite horizon loss
14:  $\hat{G}(x^{(s:H+s)})\leftarrow\hat{G}(x^{(s:H+s)})+\gamma^{H}\hat{g}^{p}(\varphi(x^{(H)},s{+}H))$ $\triangleright$ approx. terminal loss
15:  end for
16:  $x^{(1:H)}_{\star}\leftarrow\arg\max_{x^{(1:H)}\in\mathcal{C}}\hat{G}(x^{(1:H)})$
17:  $x_{s+1}\leftarrow x_{\star}^{(1)}$ $\triangleright$ execute only the first step of the best path
18:  $x^{\prime}_{s+1}\leftarrow\varphi(x_{s+1},s+1)$
19:  $\hat{g}^{i}_{s+1}\leftarrow\mathcal{A}\left(\hat{g}^{i}_{s},x^{\prime}_{s+1},y_{s+1}\right)$ $\triangleright$ update regressor on instantaneous loss
20:  $\hat{g}^{p}_{s+1}\leftarrow\mathcal{A}\left(\hat{g}^{p}_{s},x^{\prime}_{s+1},y^{\prime}_{s+1}\right)$ $\triangleright$ update regressor on cumulative loss
21:end for
22:return $\hat{g}^{i},\hat{g}^{p},x_{s+1}$ $\triangleright$ updated regressors and states
23:

### v3

@@
Optimal control of the future is the next frontier for AI. Current approaches to this problem are typically rooted in reinforcement learning (RL). RL is mathematically distinct from supervised learning, which has been the main workhorse for the recent achievements in AI. Moreover, RL typically operates in a stationary environment with episodic resets, limiting its utility. Here, we extend supervised learning to address learning to control in non-stationary, reset-free environments. Using this framework, called ”Prospective Learning with Control” (PL+C), we prove that under certain fairly general assumptions, empirical risk minimization (ERM) asymptotically achieves the Bayes optimal policy. We then consider a specific instance of prospective learning with control, foraging—which is a canonical task for any mobile agent—be it natural or artificial. We illustrate that modern RL algorithms fail to learn in these non-stationary reset-free environments, and even with modifications, they are orders of magnitude less efficient than our prospective foraging agents.
@@ Introduction
9:for $s=1>T$ do $\triangleright$ Online updates $\hat{g}^{i}$ and $\hat{g}^{p}$ and $x_{s}$
10:    $\mathcal{C}\leftarrow\textsc{AllPaths}(x_{s},H)$ $\triangleright$ each $x^{(s:H+s)}\in\mathcal{C}$ is a length-$H$ trajectory, $(x^{(s)},\dots,x^{(H+s)})$
11:  for $x^{(s:H+s)}\in\mathcal{C}$ do
12:     $\hat{G}(x^{(s:H+s})\leftarrow\sum_{h=1}^{H}\gamma^{\,h-1}\;\mathrm{\hat{g}^{i}}(\varphi(x^{(h)},s{+}h))$ $\triangleright$ approx. finite horizon loss
13:     $\hat{G}(x^{(s:H+s)})\leftarrow\hat{G}(x^{(s:H+s)})+\gamma^{H}\hat{g}^{p}(\varphi(x^{(H)},s{+}H))$ $\triangleright$ approx. terminal loss
14:  end for
15:    $x^{(1:H)}_{\star}\leftarrow\arg\max_{x^{(1:H)}\in\mathcal{C}}\hat{G}(x^{(1:H)})$
16:    $x_{s+1}\leftarrow x_{\star}^{(1)}$ $\triangleright$ execute only the first step of the best path
17:    $x^{\prime}_{s+1}\leftarrow\varphi(x_{s+1},s+1)$
18:    $\hat{g}^{i}_{s+1}\leftarrow\mathcal{A}\left(\hat{g}^{i}_{s},x^{\prime}_{s+1},y_{s+1}\right)$ $\triangleright$ update regressor on instantaneous loss
19:    $\hat{g}^{p}_{s+1}\leftarrow\mathcal{A}\left(\hat{g}^{p}_{s},x^{\prime}_{s+1},y^{\prime}_{s+1}\right)$ $\triangleright$ update regressor on cumulative loss
20:end for